The equilibrium constant, K. for a neutralization
Answer #1

Let's write the equations for Kn, Ka, Kb and Kw.

K_{n}=\frac{[NH_{4}^{+}][COOH^{-}]}{[NH_{3}][HCOOH]}

K_{a}=\frac{[H^{+}][COOH^{-}]}{[HCOOH]}

K_{b}=\frac{[NH_{4}^{+}][OH^{-}]}{[NH_{3}][H_{2}O]}

K_{w}=\frac{[H^{+}][OH^{-}]}{[H_{2}O]}

Therefore, by substituting each of the above terms into our equation for Kn, we can get the following.

K_{n}=\frac{K_{b}K_{a}}{K_{w}}

K_{n}=\frac{1.8*10^{-5}*1.8*10^{-4}}{1*10^{-14}}=3.24*10^{5}